# Derivative-Free Guidance in Continuous and Discrete Diffusion Models with Soft Value-Based Decoding

### v3

@@ 4 Soft Value-Based Decoding in Diffusion Models
In this section, we first present the motivation behind developing our new algorithm. We then introduce our algorithm, which satisfies desired properties, i.e., the lack of need for fine-tuning or constructing differentiable models.
@@ 4.1 Key Observation
We introduce several key concepts. First, for $t\in[T+1,\cdots,1]$, we define the soft value function:
$v_{t-1}(\cdot):=\alpha\log\mathbb{E}_{x_{0}\sim p^{\mathrm{pre}}(x_{0}|x_{t-1})}\left[\exp\left(\frac{r(x_{0})}{\alpha}\right)|x_{t-1}=\cdot\right],$
where $\mathbb{E}_{\{p^{\mathrm{pre}}\}}[\cdot]$ is induced by $\{p^{\mathrm{pre}}_{t}(\cdot|x_{t+1})\}_{t=T}^{0}$. This value function represents the expected future reward at $t=0$ from the intermediate noisy state at $t-1$. Next, we define the following soft optimal policy (denoising process) $p^{\star,\alpha}_{t-1}:\mathcal{X}\to\Delta(\mathcal{X})$ weighted by value functions $v_{t-1}:\mathcal{X}\to\mathbb{R}$:
@@
Here, we refer to $v_{t}$ as soft value functions, $p^{\star,\alpha}_{t}$ as soft optimal policies, respectively, because they literally correspond to soft value functions and soft optimal policies where we embed diffusion models into entropy-regularized MDPs, as demonstrated in Uehara et al. (2024).
@@
While Uehara et al. (2024) presents this theorem, they use it primarily to interpret their RL-based fine-tuning introduced in Fan et al. (2023); Black et al. (2023). In contrast, our work explores how to convert this into a new fine-tuning-free optimization algorithm.
@@
Theorem 1 states that if we can hypothetically sample from $\{p^{\star,\alpha}_{t}\}_{t=T}^{0}$, we can sample from the target distribution $p^{(\alpha)}$. However, there are two challenges in sampling from each $p^{\star,\alpha}_{t-1}$:
  1. 1.
    The soft-value function $v_{t-1}$ in $p^{\star,\alpha}_{t-1}$ is unknown.
  2. 2.
    It is unnormalized (i.e., calculating the normalizing constant is often hard).
We address the first challenge later in Section 4.3. Assuming the first challenge is resolved, we consider how to tackle the second challenge. A natural approach is to use importance sampling (IS):
$p^{\star,\alpha}_{t-1}(\cdot|x_{t},c)\approx\sum_{m=1}^{M}\frac{w^{\langle m\rangle}_{t-1}}{\sum_{j}w^{\langle j\rangle}_{t-1}}\delta_{x^{\langle m\rangle}_{t-1}},\quad\{x^{\langle m\rangle}_{t-1}\}_{m=1}^{M}\sim p^{\mathrm{pre}}_{t-1}(\cdot\mid x_{t}),$
@@
Difference over Best-of-N.
Best-of-N, which generates multiple samples and selects the highest reward sample, is technically considered IS where the proposal distribution is the entire $p^{\mathrm{pre}}(x_{0})=\int\prod_{t}\{p^{\mathrm{pre}}_{t}(x_{t-1}\mid x_{t})\}dx_{1:T}$. However, the use of importance sampling in our algorithm differs significantly, as we apply it at each time step to approximate each soft-optimal policy.
@@ 4.2 Inference-Time Algorithm
Now, we introduce our algorithm, which leverages the observation, in Algorithm 1. Our algorithm is an iterative sampling method that integrates soft value functions into the standard inference procedure of pre-trained diffusion models. Each step is designed to approximately sample from a value-weighted policy $\{p^{\star,\alpha}_{t}\}_{t=T}^{0}$.
We have several important things to note.
  • When $\alpha=0$, Line 4 corresponds to $\zeta_{t-1}=\mathop{\mathrm{argmax}}_{m\in[1,\cdots,M]}\hat{v}_{t-1}(x^{\langle m\rangle}_{t-1}).$ In practice, we often recommend this choice.
  • A typical choice we recommend is $M=[5,\cdots,20]$. The performance with varying $M$ values will be discussed in Section 6.
  • Line 3 can be computed in parallel at the expense of additional memory.
  • When the normalizing constant can be calculated relatively easily (e.g., in discrete diffusion with small $K,L$), we can directly sample from $\{p^{\star,\alpha}_{t}\}_{t=T}^{0}$.
  • We can also use a proposal distribution different from $p^{\mathrm{pre}}_{t-1}$. For the extension, refer to Section C.
@@ 4.3 Learning Soft Value Functions
Next, we describe how to learn soft value functions $v_{t}(x)$ in practice. We propose two main approaches: the Monte Carlo regression approach and the posterior mean approximation approach.
@@
$v_{t}(x_{t}):=\alpha\log\mathbb{E}_{x_{0}\sim p^{\mathrm{pre}}(x_{0}|x_{t})}[\exp(r(x_{0})/\alpha)|x_{t}]\approx\log(\exp(\mathbb{E}_{x_{0}\sim p^{\mathrm{pre}}(x_{0}|x_{t})}[r(x_{0})|x_{t}]))=v^{\prime}_{t}(x_{t}).$ (1)
@@
Note that technically, without approximation introduced in (1), we can estimate $v_{t}$ by regressing $\exp(r(x_{0})/\alpha)$ onto $x_{t}$ based on the original definition. This approach may work in many cases. However, when $\alpha$ is very small, the scaling of $\exp(r(\cdot)/\alpha)$ tends to be excessively large. Due to this concern, we generally recommend using using Algorithm 2.
Remark 2 (Another way of constructing datasets).
We have used the backward process in pre-trained diffusion models to collect training datasets in Algorithm 2. Given that the backward and forward processes are designed to match marginally during the training of diffusion models as discussed in Section 3.1, we can also use the forward process in pre-trained diffusion models to construct datasets.
@@
Technically, another method for learning value functions is available such as soft-Q-learning (Section D), by leveraging soft-Bellman equations in diffusion models Uehara et al. (2024, Section 3) (as known in standard RL contexts (Levine, 2018; Haarnoja et al., 2017; Geist et al., 2019)). However, since we find Monte Carlo approaches to be more stable, we recommend them over soft-Q-learning.
@@
5 Advantages, Limitations, and Extensions of SVDD
So far, we have detailed our algorithm, SVDD. In this section, we discuss its advantages, limitations, and extensions.
@@ 5.1 Advantages
Unlike classifier guidance, SVDD  does not require differentiable proxy models, as there is no need for derivative computations. For example, if $r$ is non-differentiable feedback (e.g., physically-based simulations, QED, SA in molecule generation), our method SVDD-PM can directly utilize such feedback without constructing differentiable proxy models. In cases where non-differentiable feedback is costly to obtain, proxy reward models may still be required, but they do not need to be differentiable; thus, non-differentiable features or non-differentiable models based on scientific knowledge (e.g., molecule fingerprints, GNNs) can be leveraged. Similarly, when using SVDD-MC, while a value function model is required, it does not need to be differentiable, unlike classifier guidance.
Additionally, compared to approaches that involve derivatives (like classifier guidance or DPS), our algorithm may be more memory efficient at inference time, particularly when $M$ is moderate, and can be directly applied to discrete diffusion models mentioned in Example 2.
Proximity to pre-trained models (robust to reward over-optimization).
Since samples are consistently generated from pre-trained diffusion policies at each step, they are ensured to remain within the natural space defined by pre-trained diffusion models.
This is especially advantageous when rewards to be optimized by SVDD are learned from offline data. In such cases, learned reward functions may be inaccurate for out-of-distribution samples (Uehara et al., 2024). Consequently, conventional fine-tuning methods often suffer from over-optimization (reward hacking) by exploiting these out-of-distribution regions (Fan et al., 2023; Clark et al., 2023). Given that non-natural design spaces typically encompass a significant portion of out-of-distribution regions of the offline data, maintaining proximity to pre-trained models acts as a regularization mechanism against these regions.
@@ 5.2 Potential Limitations
Our approach requires more computational resources (if not parallelized) or memory (if parallelized), specifically (roughly) $M$ times more than standard inference methods. Considering this aspect, we conduct a fair comparison between SVDD and best-of-N in Section 6.
@@
The proximity to pre-trained models might be a disadvantage if significant changes to the pre-trained models are desired. We acknowledge that RL-based fine-tuning could be more effective for this purpose than our algorithm.
5.3 Extensions
Using a likelihood/classifier as a reward.
While we primarily consider scenarios where reward models are regression models, by adopting a similar strategy in (Zhao et al., 2024), they can be readily replaced with classifiers or likelihood functions in the context of solving inverse problems (Chung et al., 2022; Bansal et al., 2023).
Application to fine-tuning.
Our SVDD, can also be naturally extended to fine-tuning by generating samples with SVDD, and then using these samples for supervised fine-tuning.
5.4 Comparison between SVDD and Standard SMC-Methods
The algorithm most closely related to SVDD, are several SMC-Methods for guidance (Wu et al., 2024; Trippe et al., 2022; Dou and Song, 2024; Phillips et al., 2024; Cardoso et al., 2023), which utilizes “standard” sequential Monte Carlo (Del Moral and Doucet, 2014; Kitagawa, 1993) in diffusion models. Although originally designed for conditioning, it can technically be applied to reward maximization. However, they require resampling across the “entire” batch, complicating parallelization. Additionally, when batch sizes are small, as is often the case in recent large diffusion models, its performance may be suboptimal. In contrast, our method is easily parallelizable and can achieve strong performance even with a batch size of $1$, as resampling occurs for each individual sample. For further details, see Section A.
@@ 6.1 Settings
We compare the following methods.
@@
  • Best-of-N (Nakano et al., 2021): We generate samples from pre-trained models and select the top $1/N$ samples. This selection is made to ensure that the computational time during inference is approximately equivalent to that of our proposal.
  • DPS (Chung et al., 2022): DPS is a widely used enhancement of classifier guidance. Although the original work was not designed for discrete diffusion, we employ specific heuristics to adapt it for this purpose (Section E).
  • SVDD (Ours): We implement SVDD-MC and SVDD-PM. We generally set $M=20$ for images and $M=10$ for other domains, additionally we set $\alpha=0$.
@@
We provide details on the pre-trained diffusion models and downstream reward functions used. For further information, refer to Section E.
  • Images: We use Stable Diffusion v1.5 as the pre-trained diffusion model ($T=50$). For downstream reward functions, we use compressibility and aesthetic scores (LAION Aesthetic Predictor V2 in Schuhmann (2022)), as employed by Black et al. (2023); Fan et al. (2023). Note that compressibility is non-differentiable reward feedback.
  • Molecules: We use GDSS (Jo et al., 2022), trained on ZINC-250k (Irwin and Shoichet, 2005), as the pre-trained diffusion model ($T=1000$). For downstream reward functions, we use QED and SA calculated by RDKit, which are non-differentiable feedback. Here, we renormalize SA to $(10-\mathrm{SA})/9$ so that a higher value indicates better performance.
  • DNAs (Enhancers): We use the discrete diffusion model (Sahoo et al., 2024), trained on datasets from Gosai et al. (2023), as our pre-trained diffusion model ($T=128$). For the downstream reward function, we use an Enformer model (Avsec et al., 2021) to predict activity in the HepG2 cell line.
  • RNAs (5’UTRs): We use the discrete diffusion model (Sahoo et al., 2024) as the pre-trained diffusion model ($T=128$). For downstream reward functions, we employ a reward model that predicts the mean ribosomal load (MRL) measured by polysome profiling (trained on datasets from Sample et al. (2019)) and stability (trained on datasets from (Agarwal and Kelley, 2022)).
@@ 6.2 Results
(a) Images: compressibility
(b) Images: aesthetic score
(c) Molecules: QED
(d) Molecules: SA
(e) Enhancers
(f) 5’UTRs: MRL
(g) 5’UTRs: stability
Figure 2: We show the histogram of generated samples in terms of reward functions.
@@
Figure 3: We show generated samples from our proposal. For more samples, refer to Section E.2.
We compare the baselines with our two proposed methods. The performance is shown in Figure 2, and the generated samples are presented in Figure 3.
Overall, our proposal outperforms the baseline methods (Best-of-N and DPS), as evidenced by higher rewards for samples generated in large quantities. In Section E, we compare the top $10^{\text{th}}$ and $20^{\text{th}}$ quantiles from each algorithm and confirm that SVDD always outperforms the baselines. This indicates that our algorithm can generate high-reward samples that Best-of-N and DPS cannot.
  • Compared to Best-of-N, although the rewards for generating samples in smaller quantities could be lower with our algorithm, this is expected because our algorithm generates samples with high likelihood $p^{\mathrm{pre}}(x|c)$ in pre-trained diffusion models, but with possibly lower rewards.
  • Compared to DPS, our algorithm consistently achieves better performance. Notably, in molecular generation, DPS is highly ineffective due to the non-differentiable nature of the original feedback.
The relative performance of our two proposed methods (SVDD-MC or SVDD-PM) appears to be domain-dependent. Generally, SVDD-PM may be more robust since it does not require additional learning (i.e., it directly utilizes reward feedback). The performance of SVDD-MC depends on the success of value function learning, which is discussed in Section E.
Figure 4: Performance of our algorithm SVDD as $M$ varies, for image generation while optimizing the aesthetic score.
Ablation studies in terms of $M$.
We plot the performance as $M$ varies. The performance gradually reaches a plateau as $M$ increases. This tendency is seen in all other domains.

### v4

@@ 4 Soft Value-Based Decoding in Diffusion Models
First, we present the motivation behind developing our new algorithm. We then introduce SVDD, which satisfies our desired properties, i.e., the lack of need for fine-tuning or constructing differentiable models.
@@ 4.1 Key Observation
We introduce several key concepts. First, we define the soft value function:
$\textstyle t\in[T+1,\cdots,1];v_{t-1}(\cdot):=\alpha\log\mathbb{E}_{x_{0}\sim p^{\mathrm{pre}}(x_{0}|x_{t-1})}\left[\exp\left(\frac{r(x_{0})}{\alpha}\right)|x_{t-1}=\cdot\right],$
where $\mathbb{E}_{\{p^{\mathrm{pre}}\}}[\cdot]$ is induced by $\{p^{\mathrm{pre}}_{t}(\cdot|x_{t+1})\}_{t=T}^{0}$. This value function represents the expected future reward at $t=0$ from the intermediate noisy state at $t-1$.
Next, we define the following soft optimal policy (denoising process) $p^{\star,\alpha}_{t-1}:\mathcal{X}\to\Delta(\mathcal{X})$ weighted by value functions $v_{t-1}:\mathcal{X}\to\mathbb{R}$:
@@
Here, $v_{t}$ are soft value functions and $p^{\star,\alpha}_{t}$ are soft optimal policies, because they literally correspond, respectively, to soft value functions and soft optimal policies, where we embed diffusion models into entropy-regularized MDPs (Geist et al., 2019), as demonstrated in Uehara et al. (2024).
@@
While Uehara et al. (2024) presents this theorem, they use it primarily to interpret RL-based fine-tuning methods in Fan et al. (2023); Black et al. (2023). In contrast, our work explores how to convert this into a new fine-tuning-free optimization algorithm.
@@
Theorem 1 states that if we can hypothetically sample from $\{p^{\star,\alpha}_{t}\}_{t=T}^{0}$, we can sample from the target distribution $p^{(\alpha)}$. However, there are two challenges in sampling from each $p^{\star,\alpha}_{t-1}$: (1) the soft-value function $v_{t-1}$ in $p^{\star,\alpha}_{t-1}$ is unknown, and (2) it is unnormalized (i.e., calculating the normalizing constant is hard).
We address the first challenge in Section 4.3. Assuming the first challenge is resolved, we consider how to tackle the second challenge. A natural approach is to use importance sampling (IS):
$p^{\star,\alpha}_{t-1}(\cdot|x_{t},c)\approx\sum_{m=1}^{M}\frac{w^{\langle m\rangle}_{t-1}}{\sum_{j=1}^{M}w^{\langle j\rangle}_{t-1}}\delta_{x^{\langle m\rangle}_{t-1}},\quad\{x^{\langle m\rangle}_{t-1}\}_{m=1}^{M}\sim p^{\mathrm{pre}}_{t-1}(\cdot\mid x_{t}),$
@@
Note that Best-of-N, which generates multiple samples and selects the highest reward sample, is technically considered IS, where the proposal distribution is the entire $p^{\mathrm{pre}}(x_{0})=\int\prod_{t}\{p^{\mathrm{pre}}_{t}(x_{t-1}\mid x_{t})\}dx_{1:T}$. However, the use of importance sampling in our algorithm differs significantly, as we apply it at each time step to approximate each soft-optimal policy.
@@ 4.2 Inference-Time Algorithm
4:     $x_{t-1}\leftarrow x^{\langle\zeta_{t-1}\rangle}_{t-1}$ after selecting an index: $\zeta_{t-1}\sim\mathrm{Categorical}\left(\left\{\frac{w^{\langle m\rangle}_{t-1}}{\sum_{j=1}^{M}w^{\langle j\rangle}_{t-1}}\right\}_{m=1}^{M}\right),\,$
@@
Now, by leveraging the observation in Algorithm 1, we introduce our algorithm. Our algorithm is an iterative sampling method that integrates soft value functions into the standard inference procedure of pre-trained diffusion models. Each step is designed to approximately sample from a value-weighted policy $\{p^{\star,\alpha}_{t}\}_{t=T}^{0}$.
We note several key points.
  • When $\alpha=0$, Line 4 corresponds to $\zeta_{t-1}=\mathop{\mathrm{argmax}}_{m\in[1,\cdots,M]}\hat{v}_{t-1}(x^{\langle m\rangle}_{t-1}).$ In practice, we often recommend this choice. This is the default choice in Section 7.
  • A typical choice we recommend for $M$ is from $5$ to $20$. The performance with varying $M$ values will be discussed in Section 7.
  • Line 3 can be computed in parallel at the cost of additional memory (scaled by $M$). If Line 3 is not computed in parallel, the computational time in SVDD would be approximately $M$ times that of the standard inference procedure. We will check it in Section 7.
  • In special cases where the normalizing constant can be calculated relatively easily (e.g., in discrete diffusion with small $K,L$), we can directly sample from $\{p^{\star,\alpha}_{t}\}_{t=T}^{0}$.
  • A proposal distribution different from $p^{\mathrm{pre}}_{t-1}$ in line 3 can be applied (see Section D). For instance, classifier guidance or its variants may be used to obtain better proposal distributions than those from the pure pre-trained model.
@@ 4.3 Learning Soft Value Functions
Next, we describe how to obtain soft value functions $v_{t}(x)$ in practice. We propose two main approaches: a Monte Carlo regression approach and a posterior mean approximation approach.
@@
$v_{t}(x_{t})=\alpha\log\mathbb{E}_{x_{0}\sim p^{\mathrm{pre}}(\cdot|x_{t})}[\exp(r(x_{0})/{\alpha}|x_{t}]\approx\log(\exp(\mathbb{E}_{x_{0}\sim p^{\mathrm{pre}}(x_{t})}[r(x_{0})|x_{t}]))=v^{\prime}_{t}(x_{t}).$ (1)
@@
Note that technically, without the approximation introduced in (1), we can estimate $v_{t}$ by regressing $\exp(r(x_{0})/\alpha)$ onto $x_{t}$ based on the original definition. This approach may work in many cases. However, when $\alpha$ is very small, the scaling of $\exp(r(\cdot)/\alpha)$ tends to be excessively large. Due to this concern, we generally recommend using Algorithm 2.
@@
Technically, another method for learning value functions is available such as soft-Q-learning (Section E), by leveraging soft-Bellman equations in diffusion models Uehara et al. (2024, Section 3) However, since we find Monte Carlo approaches to be more stable, we recommend them over soft-Q-learning.
@@
5 Advantages, Extensions, Limitations of SVDD
We discuss the advantages, extensions, and limitations of SVDD.
@@ 5.1 Advantages
Unlike classifier guidance, SVDD  does not require differentiable proxy models, as there is no need for derivative computations. For example, if $r$ is non-differentiable feedback (e.g., physically-based simulations for docking scores in molecule generation), our method SVDD-PM can directly utilize such feedback without constructing differentiable proxy models. In cases where non-differentiable computational feedback is costly to obtain, the usage of proxy reward models may still be preferred, but they do not need to be differentiable; thus, non-differentiable features or non-differentiable models based on scientific knowledge (e.g., molecule fingerprints, GNNs) can be leveraged. Similarly, when using SVDD-MC, while a value function model is required, it does not need to be differentiable, unlike classifier guidance. Additionally, compared to approaches that involve derivatives (like classifier guidance or DPS), SVDD can be directly applied to discrete diffusion models mentioned in Example 2.
5.2 Extensions
Using a likelihood/classifier as a reward.
While we primarily consider scenarios where reward models are regression models, by adopting a similar strategy to that in Zhao et al. (2024), they can be readily replaced with classifiers or likelihood functions in the context of solving inverse problems or conditioning (Chung et al., 2022; Bansal et al., 2023).
Fine-tuning by distilling SVDD.
The inference-time cost may become slow as $M$ increases in SVDD,. This issue can be mitigated by policy distillation, that is further fine-tuning diffusion models to align them closely with policies from SVDD(Salimans and Ho, 2022; Kim et al., 2023). We leave this aspect for future work.
@@ 5.3 Potential Limitations
Our approach requires more computational resources (if not parallelized) or memory (if parallelized), approximately $M$ times more than standard inference methods, as noted in Section 4.2. Taking this aspect into account, we compare SVDD, with baselines such as best-of-N in our experimental section (Section 7). For gradient-based approaches like classifier guidance and DPS, while a direct comparison with SVDD is challenging, it is important to note that these methods also incur additional computational and memory complexity due to the backward pass, which SVDD avoids. Lastly, it is important to note that this additional inference-time burden can be alleviated through distillation, as discussed in Section 5.2.
@@
If significant changes to the pre-trained models are desired, we acknowledge that RL-based fine-tuning (Black et al., 2023; Fan et al., 2023) could be more effective than SVDD for this purpose in certain scenarios, such as image examples. However, this proximity to pre-trained models could also be advantageous in the sense that it is robust against reward optimization, which conventional fine-tuning methods often suffer from by exploiting these out-of-distribution regions (Uehara et al., 2024). Lastly, in cases where reward backpropagation (Prabhudesai et al., 2023; Clark et al., 2023; Uehara et al., 2024) is not applicable, particularly in scientific domains for RL-based fine-tuning, we may need to rely on PPO. However, PPO is often mode-seeking and unstable, highlighting the challenges of RL-based fine-tuning in certain scenarios.
6 Comparison between SVDD and SMC-Based Methods
SMC-based methods for diffusion models are closely related to SVDD. These approaches (Wu et al., 2024; Trippe et al., 2022; Dou and Song, 2024; Phillips et al., 2024; Cardoso et al., 2023) use SMC (Del Moral and Doucet, 2014) for sampling from diffusion models. While they are originally designed for conditioning (by setting rewards as classifiers), they can also be applied to reward maximization. Notably, similar to our work, these methods do not require differentiable models.
However, these SMC methods are not tailored to reward maximization. Most importantly, they involve resampling across the “entire” batch, which complicates parallelization. Additionally, when batch sizes are small, as is often the case with recent large diffusion models, performance may be suboptimal, since the SMC theoretical guarantees hold primarily with large batch sizes. Even with larger batch sizes, using SMC for reward maximization can result in a loss of diversity across the entire batch, since the effective sample size based on weights, which is a standard diversity measure in SMC, does not ensure “real” diversity in the generated samples. In contrast, our method is highly parallelizable, performs well even with small batch sizes (as low as $1$), and maintains diversity with larger batch sizes, as sampling is conducted on a “per-sample basis” (Line 4). We empirically validate this in Section 7. We provide further details and experiments in Appendix B.
@@ 7.1 Settings
We compare SVDD to several representative methods capable of performing reward maximization during inference, discussed in Section 2.
@@
  • Best-of-N: We generate samples from pre-trained models and select the top $1/N$ samples. This selection is made to ensure that the computational time during inference is approximately equivalent to that of SVDD.
  • DPS (Chung et al., 2022): It is a widely used training-free version of classifier guidance. For discrete diffusion, we combine it with the state-of-the-art approach (Nisonoff et al., 2024).
  • SMC-Based Methods (SMC): Methods discussed in Section 6 and Appendix B, which do not require differentiable models, like SVDD.
  • SVDD (Ours): We implement SVDD-MC and SVDD-PM. We generally set $M=20$ for images and $M=10$ for other domains, and $\alpha=0$. Recall $M$ is the duplication size in the IS part.
@@
We provide details on the pre-trained diffusion models and downstream reward functions used. For further information, refer to Section F.
  • Images: We use Stable Diffusion v1.5 as the pre-trained diffusion model ($T=50$). For downstream reward functions, we use compressibility and aesthetic scores (LAION Aesthetic Predictor V2 in Schuhmann (2022)), as employed by Black et al. (2023); Fan et al. (2023). Compressibility is a non-differentiable reward feedback.
  • Molecules: We use GDSS (Jo et al., 2022), trained on ZINC-250k (Irwin and Shoichet, 2005), as the pre-trained diffusion model ($T=1000$). For downstream reward functions, we use drug-likeness (QED) and synthetic accessibility (SA) calculated by RDKit, as well as docking score (DS) calculated by QuickVina 2 (Alhossary et al., 2015), which are all non-differentiable feedback. Here, we renormalize SA to $(10-\mathrm{SA})/9$ and docking score to $\max(-\mathrm{DS},0)$, so that a higher value indicates better performance. The docking scores measure binding affinity regarding four target proteins: Parp1, 5ht1b, Jak2, and Braf following Yang et al. (2021). These tasks are critical for drug discovery.
  • DNAs (enhancers) and RNAs (5’Untranslated regions (UTRs)): We use the discrete diffusion model (Sahoo et al., 2024), trained on datasets from Gosai et al. (2023) for enhancers, and from Sample et al. (2019) for 5’UTRs, as our pre-trained diffusion model ($T=128$). For the reward functions, we use an Enformer model (Avsec et al., 2021) to predict activity of enhancers in the HepG2 cell line, and a ConvGRU model that predicts the mean ribosomal load (MRL) of 5’UTRs measured by polysome profiling, respectively (Sample et al., 2019). These tasks are highly relevant for cell and RNA therapies, respectively (Taskiran et al., 2024; Castillo-Hair and Seelig, 2021).
@@ 7.2 Results
Table 2: Top 10 and 50 quantiles of the generated samples for each algorithm (with 95% confidence intervals). Higher is better. SVDD consistently outperforms the baseline methods.
Domain Quantile Pre-Train Best-N DPS SMC SVDD-MC SVDD-PM
Image: Compress 50% -101.4 $\pm$ 0.22 -71.2 $\pm$ 0.46 -60.1 $\pm$ 0.44 -59.7 $\pm$ 0.4 -54.3 $\pm$ 0.33 -51.1 $\pm$ 0.38
10% -78.6 $\pm$ 0.13 -57.3 $\pm$ 0.28 -61.2 $\pm$ 0.28 -49.9 $\pm$ 0.24 -40.4 $\pm$ 0.2 -38.8 $\pm$ 0.23
Image: Aesthetic 50% 5.62 $\pm$ 0.003 6.11 $\pm$ 0.007 5.61 $\pm$ 0.009 6.02 $\pm$ 0.004 5.70 $\pm$ 0.008 6.14 $\pm$ 0.007
10% 5.98 $\pm$ 0.002 6.34 $\pm$ 0.004 6.00 $\pm$ 0.005 6.28 $\pm$ 0.003 6.05 $\pm$ 0.005 6.47 $\pm$ 0.004
Molecule: QED 50% 0.656 $\pm$ 0.008 0.835 $\pm$ 0.009 0.679 $\pm$ 0.024 0.667 $\pm$ 0.016 0.852 $\pm$ 0.011 0.848 $\pm$ 0.014
10% 0.812 $\pm$ 0.005 0.902 $\pm$ 0.006 0.842 $\pm$ 0.014 0.722 $\pm$ 0.009 0.925 $\pm$ 0.007 0.928 $\pm$ 0.008
Molecule: SA 50% 0.652 $\pm$ 0.007 0.834 $\pm$ 0.014 0.693 $\pm$ 0.022 0.786 $\pm$ 0.004 0.935 $\pm$ 0.010 0.925 $\pm$ 0.016
10% 0.803 $\pm$ 0.004 0.941 $\pm$ 0.008 0.844 $\pm$ 0.013 0.796 $\pm$ 0.003 1.000 $\pm$ 0.006 1.000 $\pm$ 0.010
Molecule: Docking parp1 50% 7.15 $\pm$ 0.52 10.00 $\pm$ 0.17 7.35 $\pm$ 0.43 6.90 $\pm$ 0.60 12.00 $\pm$ 0.26 11.40 $\pm$ 0.22
10% 8.59 $\pm$ 0.31 10.67 $\pm$ 0.10 9.31 $\pm$ 0.26 9.37 $\pm$ 0.36 13.25 $\pm$ 0.16 12.41 $\pm$ 0.13
Molecule: Docking 5ht1b 50% 7.20 $\pm$ 0.53 9.65 $\pm$ 0.17 7.30 $\pm$ 0.48 6.80 $\pm$ 0.35 10.50 $\pm$ 0.46 10.50 $\pm$ 0.53
10% 8.69 $\pm$ 0.32 10.28 $\pm$ 0.10 9.21 $\pm$ 0.29 9.00 $\pm$ 0.21 12.87 $\pm$ 0.28 12.30 $\pm$ 0.32
Molecule: Docking jak2 50% 7.05 $\pm$ 0.45 8.85 $\pm$ 0.17 7.30 $\pm$ 0.43 6.80 $\pm$ 0.46 10.65 $\pm$ 0.35 10.30 $\pm$ 0.45
10% 8.20 $\pm$ 0.27 9.59 $\pm$ 0.10 8.70 $\pm$ 0.26 10.00 $\pm$ 0.28 11.80 $\pm$ 0.21 11.91 $\pm$ 0.27
Molecule: Docking braf 50% 7.20 $\pm$ 0.40 9.20 $\pm$ 0.11 7.50 $\pm$ 0.23 6.90 $\pm$ 0.46 10.00 $\pm$ 0.37 9.65 $\pm$ 0.33
10% 8.59 $\pm$ 0.24 10.29 $\pm$ 0.07 9.20 $\pm$ 0.14 8.74 $\pm$ 0.28 11.30 $\pm$ 0.22 11.40 $\pm$ 0.20
Enhancers 50% 0.121 $\pm$ 0.033 1.807 $\pm$ 0.214 3.782 $\pm$ 0.299 $4.28\pm 0.02$ 5.074 $\pm$ 0.096 5.353 $\pm$ 0.231
10% 1.396 $\pm$ 0.020 3.449 $\pm$ 0.128 4.879 $\pm$ 0.179 $5.95\pm 0.01$ 5.639 $\pm$ 0.057 6.980 $\pm$ 0.138
5’UTR 50% 0.406 $\pm$ 0.028 0.912 $\pm$ 0.023 0.426 $\pm$ 0.073 $0.76\pm 0.02$ 1.042 $\pm$ 0.008 1.214 $\pm$ 0.016
10% 0.869 $\pm$ 0.017 1.064 $\pm$ 0.014 0.981 $\pm$ 0.044 $0.91\pm 0.01$ 1.117 $\pm$ 0.005 1.383 $\pm$ 0.010
@@
(e) Molecules (by SVDD): high docking scores to Parp1
(f) Molecules (by SVDD): high docking scores to 5ht1b
(g) Molecules (by SVDD): high docking scores to Jak2
(h) Molecules (by SVDD): high docking scores to Braf
Figure 2: Generated samples from SVDD. For more samples, please refer to Section F.3. Note that the surfaces and ribbons in (e)-(h) (such as the green objects in (e)) are representations of the target proteins, while the generated small molecules are displayed in the center.
We compare the baselines with our two proposed methods regarding rewards $r$ (Table 2). To demonstrate the generated samples’ validity (i.e., naturalness), we present several examples in Figure 2. In Section F.3, we also include metrics for the validity of samples.
Overall, SVDD outperformed the baseline methods (Best-of-N, SMC, DPS, SMC), as evidenced by higher quantiles. Furthermore, for both molecules and images, the samples generated by SVDD were valid. This suggests that SVDD can generate high-reward valid samples that Best-of-N, DPS, and SMC often struggle to generate or, in some cases, nearly fail to do.
The relative performance of our SVDD-MC and SVDD-PM appears to be domain-dependent. Generally, SVDD-PM may be more robust since it does not require additional learning (i.e., it directly utilizes reward feedback). The performance of SVDD-MC depends on the success of value function learning discussed in Section F.
(a) Performance of SVDD as $M$ varies for image generation while optimizing aesthetic score.
(b) Performance of SVDD as $M$ varies for molecule generation while optimizing the QED score.
(c) GPU time and max memory of SVDD as $M$ varies for image generation (aesthetic scores).
(d) GPU time and max memory of SVDD as $M$ varies for molecule generation (QED).
Figure 3: Ablation studies with respect to $M$ for SVDD. Note 3(c) and 3(d) indicate that the computational time does not scale linearly with $M$, whereas memory usage scales linearly.
Ablation studies in terms of the duplication size $M$.
We assessed the performance of SVDD-PM (when calculating value functions in Line 3 in a non-parallel manner) along with the computational and memory complexity as $M$ varies. First, across all domains, the performance gradually plateaus as $M$ increases (Figure 3(a) and 3(b)). Second, computational complexity increases linearly with $M$, while memory complexity remains nearly constant (Figure 3(c) and 3(d)). This behavior is expected, as previously noted in Section 4.2. The comparison with Best-of-N in Table 2 is made with this consideration in mind.